# Astrometry-Only Detection of Microlensing Events with Gaia

### v1

@@ 1 Introduction
Photometric microlensing has been extensively studied by decades-long surveys such as the Optical Gravitational Lensing Experiment (OGLE; Udalski 2003) or Microlensing Observations in Astrophysics (MOA; Bond et al. 2001). However, only recently have technological advances and high-resolution observatories enabled the observation of astrometric microlensing (Dominik & Sahu 2000; Belokurov & Evans 2002) and allowed the detection and characterisation of objects that are difficult to observe through photometric methods alone, such as low-mass stars, white dwarfs, neutron stars, and black holes (Sahu et al. 2017; Zurlo et al. 2018; Dong et al. 2019; Sahu et al. 2022; Jabłońska, M. et al. 2022).
@@
The Gaia mission (Gaia Collaboration et al. 2016), launched by the European Space Agency, has revolutionised astrometry by providing high-precision measurements of the positions, motions and parallaxes of more than a billion stars (Gaia Collaboration et al. 2018, 2021, 2023). Gaia’s precise astrometric data are particularly suited for identifying microlensing events. Belokurov & Evans (2002) estimated that approximately 25,000 stars will exhibit noticeable centroid shifts during the mission. Gaia results are published in Data Releases (DRs), which can be used to predict, find, or confirm astrometric microlensing events. Specifically, Klüter, J. et al. (2018); Klüter et al. (2022) and Su et al. (2023) used Gaia data to predict between 4,500 and 5,672 astrometric microlensing events from 2010 to 2070.
@@
Despite the high number of expected astrometric microlensing events detected by Gaia, confirmed observations remain limited to a single candidate. This candidate was identified by Jabłońska, M. et al. (2022) among the 363 photometric microlensing events in Gaia DR3 (Wyrzykowski, Ł. et al. 2023). With the upcoming Gaia Data Release 4 in 2026, which will provide the complete astrometric time-series for all stars, there is a growing requirement for software tools to filter and detect astrometric microlensing events. By not relying on photometric data, such software would allow us to detect and study a larger sample of microlensing events, thereby enhancing our understanding of lensing objects. Chen et al. (2023) developed a data analysis pipeline for detecting astrometric microlensing events caused by black holes, with an emphasis on obtaining dark matter constraints. Their approach focuses on the detection of black holes using astrometry-only data of Gaia DR4, while in our work we generalise the approach to other types of remnants.
@@
In this paper, we present the development and application of the Gaia Astrometric Microlensing Events (GAME) Filter. This software tool is designed to identify microlensing events within Gaia’s dataset and derive the properties of the lensing objects. We tested and validated the performance of GAME Filter using mock Gaia observations across a range of magnitudes, numbers of observations, and durations of events. Additionally, we assessed the rate of false positive detections involving binary systems misclassified as microlensing events. As it will be shown in this paper, GAME Filter successfully recovers microlensing parameters. This indicates that GAME Filter holds great potential for providing insights into microlensing events and their associated lensing objects.
@@
The paper is structured as follows. In Section 2, we describe the basics of astrometric microlensing, the process of generating mock Gaia observations, and the main properties of GAME filter. In Section 3, we present the results, focussing on the performance of GAME Filter under various conditions. Section 4 provides a discussion, and Section 5 our main findings.
@@ 2 Methodology
In this section, we present the main astrometric microlensing relations. We describe how we generated mock Gaia observations with astromet [^1] https://github.com/zpenoyre/astromet.py and the main functionalities of GAME filter adapted for jaxtromet [^2] https://github.com/maja-jablonska/jaxtromet.py. astromet and jaxtromet are Python packages developed to simulate astrometric tracks of single sources, microlensing events, and binary systems (e.g., Penoyre et al. 2022). astromet is optimised for parallel CPU processing, while jaxtromet leverages the Python library JAX [^3] https://github.com/google/jax to perform calculations on GPUs, enhancing computational efficiency. Simulating 50,000 mock Gaia observations requires approximately 8 hours while filtering and identifying microlensing events takes an additional 13 hours on a CPU with 24 cores.
@@ 2.1 Astrometric microlensing
The separation between the source and the lens, normalized by $\theta_{\mathrm{E}}$, is given by
$u(t)=\sqrt{\left(\frac{t-t_{0}}{t_{\mathrm{E}}}+\delta t\right)^{2}+\left(u_{0}+\delta\beta\right)^{2}},$ (3)
where $u_{0}$ is the impact parameter — the closest approach between the lens and the source which occurs at time $t_{0}$. The displacement vector $(\delta\tau,\delta\beta)$ due to parallax is
$(\delta\tau,\delta\beta)=\boldsymbol{\pi}_{\mathrm{E}}\Delta\mathbf{s}=(\boldsymbol{\pi}_{\mathrm{E}}\cdot\Delta\mathbf{s},\boldsymbol{\pi}_{\mathrm{E}}\times\Delta\mathbf{s}),$ (4)
where $\Delta\mathbf{s}$ is the positional offset of the Sun in the geocentric frame (Gould 2004). The microlensing parallax $\boldsymbol{\pi}_{\mathrm{E}}$ is parallel to the proper motion vector and can be decomposed into equatorial components east $\pi_{\mathrm{EE}}$ and north $\pi_{\mathrm{EN}}$. This, in turn, can be used to determine the distance to the lens
@@
where $\pi_{\mathrm{E}}=|\boldsymbol{\pi}_{\mathrm{E}}|=\sqrt{\pi_{\mathrm{EE}}^{2}+\pi_{\mathrm{EN}}^{2}}$.
@@
In general, the light from the source can blend with a luminous lens, affecting the centroid deviation described by Equation 7 (Hog et al. 1995; Miyamoto & Yoshii 1995; Walker 1995). However, in our study, we focused on dark lenses and ignored any additional contribution to the light observed from sources not related to the event. In this scenario, the shift of the light centroid traces out an ellipse in the sky (e.g., Dominik & Sahu 2000). The shape and orientation of this ellipse could in principle provide information about the lensing parameters, including the mass and distance of the lens. By fitting the theoretical model to the observed astrometric shifts, it is possible to derive the properties of the lensing object (Belokurov & Evans 2002).
@@ 2.2 Generating mock Gaia observations
We generated mock Gaia observations corresponding to Gaia DR4 run from 2014.5 to 2020 with astromet. To obtain the Gaia scanning laws, the scanning angles $\varphi_{\mathrm{obs}}$ as a function of the observation time $t_{\mathrm{obs}}$, we used the Gaia Observing Schedule Tool (GOST) [^4] https://gaia.esac.esa.int/gost/. We set the reference epoch at $t_{\mathrm{ref}}=2017.5$ and assumed that there is no blending in the observations (neither from the lens nor from unrelated sources). The position and motion of a source on the celestial sphere at $t_{\mathrm{ref}}$ is defined by its Right Ascension $\alpha_{0}$, Declination $\delta_{0}$, proper motion in Right Ascension $\mu_{\alpha^{*}}$, [^5] We used a “*” symbol for motions corrected for $\cos\delta$; $\mu_{\alpha^{*}}=\mu_{\alpha}\cos\delta$. proper motion in Declination $\mu_{\delta}$, and parallax $\varpi$. We generated seven microlensing datasets. Three data sets correspond to three different baseline G-band magnitudes $G_{0}=14$, 16.5, and 19 mag, assuming the source position $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ})$. Three data sets correspond to three different source positions $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ}),\,(277.8^{\circ},-68.3^{\circ}),\,(287.9^{\circ},45.9^{\circ})$ as shown in Table 1, assuming $G_{0}=14$, which are characterized by the number of Gaia visits $N_{\mathrm{v}}=281$, 209, and 91, respectively. [^6] These $N_{\mathrm{v}}$ values correspond to regions most frequently visited by Gaia (281 visits), regions with an approximately average visit number (91 visits), and intermediate regions (209 visits). We determine $N_{\mathrm{v}}$ using the scanninglaw Python package (https://github.com/gaiaverse/scanninglaw), as this information is not available in GOST. However, since this package does not provide data for Gaia DR4, we used $N_{\mathrm{v}}$ for Gaia DR3.. We also generate one additional data set with $t_{0}\in[2010,2014]\,$yr, $G_{0}=14$, $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ})$, meaning that the peak of the microlensing signal was located outside of the DR4 time span. To generate mock Gaia observations of microlensing events, we used five single source parameters and six microlensing parameters. The single source parameters are: $\alpha_{0}$, $\delta_{0}$, $\mu_{\alpha^{*}}$, $\mu_{\delta}$, and $\varpi$. The microlensing parameters are $u_{0}$, $\theta_{\mathrm{E}}$, $t_{0}$, $t_{\mathrm{E}}$, $\pi_{\mathrm{EE}}$, and $\pi_{\mathrm{EN}}$. The specific ranges of microlensing parameters that we used to generate mock data are shown in Table 1. For each dataset, we simulate 50,000 microlensing events.
Table 1: Microlensing parameter ranges used to generate mock Gaia astrometric observations. We consider three different sky positions observed with Gaia, while the other parameters are generated from random uniform distributions within the specified ranges.
@@
Other types of astrometric events can produce light centroid tracks similar to those of microlensing events, which can lead to event misclassification. To assess the rate of such false positive microlensing detections, we generated a mock Gaia dataset of 50,000 binary system events using a procedure similar to that used for microlensing events. The required parameters are five single source parameters and nine binary system parameters. The parameters of binary systems are the orbital period $P$, the orbital semimajor axis $a$, the orbital eccentricity $e$, the ratio between masses of both stars $q$, the ratio between luminosities of both stars $l$, the polar viewing angle $\theta_{\mathrm{v}}$, the azimuthal viewing angle $\phi_{\mathrm{v}}$, the planar projection angle $\omega_{\mathrm{v}}$, and time of periapse passage $t_{\text{peri}}$ (Penoyre et al. 2022). The specific ranges of the binary system parameters used to generate mock data are gathered in Table 2.
Table 2: Binary system parameter ranges used to generate mock Gaia astrometric observations. We consider a specific $t_{\mathrm{peri}}$, while the other parameters are generated from random uniform distributions limited by the specific parameter range. The single source parameter ranges and distributions are the same as for microlensing events (see Table 1).
@@
2.3 GAME Filter
GAME Filter [^7] https://github.com/tajjankovic/GAME-Filter/ is a software tool developed to identify microlensing events in the Gaia dataset and obtain the parameters of the event that can help derive parameters and properties of the lensing object. The software reads $x_{\mathrm{obs}}$, $x_{\mathrm{err}}$, $\Delta x_{\mathrm{obs}}$, $t_{\mathrm{obs}}$, and $\varphi_{\mathrm{obs}}$ from the Gaia data files. GAME Filter then calculates $x_{\mathrm{fit}}$, which is the deviation along $\varphi_{\mathrm{obs}}$ at $t_{\mathrm{obs}}$, for specific single source and microlensing parameters. After that, the software minimizes the renormalized microlensing unit weight error (MUWE)
a scalar parameter which indicates the goodness of the microlensing fit, where $N$ is the number of visits for a specific event. The goal of the minimization is to identify parameters that bring MUWE as close as possible to its theoretical minimum value of 1. The minimisation process uses limited memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS-B) algorithm [^8] We tested several minimisation methods and found that the L-BFGS-B method produces the most accurate results. This is mainly due to its feature of constraining individual parameters within bounds to explore the parameter space and determine the optimal single source and microlensing parameters for individual events. Appendix A provides a more detailed description of the minimisation process.
Following the minimisation process, the minimiser might stop at an incorrect local minimum, failing to find the correct solution. Consequently, we established criteria to determine when an event is recovered. These criteria are based on the value of MUWE after minimisation $\mathrm{MUWE}_{\mathrm{min}}$, L2 optimality error $L_{\mathrm{opt}}$ (see Appendix A), initial guesses, and the boundaries imposed on individual parameters. We determined the critical values of $\mathrm{MUWE}_{\mathrm{min}}$ and $L_{\mathrm{opt}}$ from their respective histograms as the values where histograms show a sharp decline. We considered an event to be recovered if the following criteria are met.
  • $1.10.9<\mathrm{MUWE}_{\mathrm{min}}<$ 1.1.
  • $L_{\mathrm{opt}}<0.015$.
  • The values of all parameters are within the boundaries imposed.
3 Results
In Table 3, we list all the simulated datasets with their abbreviated names and properties. Furthermore, we show the percentages of recovered events $P_{\mathrm{rec}}$ and the percentages of events with parameter values obtained after minimisation within 10% (20%) of their true values, denoted as $P_{10}$ ($P_{20}$). These metrics provide a direct measure of GAME filter’s accuracy. Detailed discussions on these results are provided in the following subsections.
The true parameter values used to generate mock observations are: the source’s Right Ascension ($\alpha_{0}$), Declination ($\delta_{0}$), proper motion in Right Ascension ($\mu_{\alpha^{*}}$), proper motion in Declination ($\mu_{\delta}$), and parallax ($\varpi$). The microlensing parameters are: the impact parameter ($u_{0}$), the Einstein radius ($\theta_{\mathrm{E}}$), the time of the closest approach between the source and the lens ($t_{0}$), the event timescale ($t_{\mathrm{E}}$), and the microlensing parallax components ($\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$). We refer to the parameter values obtained after minimisation as $\alpha_{0,\mathrm{min}}$, $\delta_{0,\mathrm{min}}$, $\mu_{\alpha^{*},\mathrm{min}}$, $\mu_{\delta,\mathrm{min}}$, $\varpi_{\mathrm{min}}$, $u_{\mathrm{0,min}}$, $\theta_{\mathrm{E,min}}$, $t_{\mathrm{0,min}}$, $t_{\mathrm{E,min}}$, $\pi_{\mathrm{EE,min}}$, and $\pi_{\mathrm{EN,min}}$. In our study, we omit showing the plots for $\alpha_{0,\mathrm{min}}$ and $\delta_{0,\mathrm{min}}$ because we impose tight boundaries on these parameters and they are always accurately recovered. This is motivated by the fact that the position of the observed region in the sky is always accurately known.
Table 3: Simulated dataset with their respective parameters and accuracy of parameter estimation.

### v2

@@ 1 Introduction
Photometric microlensing has been extensively studied by decades-long surveys such as the Optical Gravitational Lensing Experiment (OGLE; Udalski 2003) or Microlensing Observations in Astrophysics (MOA; Bond et al. 2001). However, only recently have technological advances and high-resolution observatories enabled the observation of astrometric microlensing (Dominik & Sahu 2000; Belokurov & Evans 2002) and allowed the detection and characterisation of objects that are difficult to observe through photometric methods alone, such as low-mass stars, white dwarfs, neutron stars, and black holes (Sahu et al. 2017; Zurlo et al. 2018; McGill et al. 2018; Dong et al. 2019; Sahu et al. 2022; Jabłońska, M. et al. 2022; Fardeen et al. 2024).
@@
The Gaia mission (Gaia Collaboration et al. 2016), launched by the European Space Agency, has revolutionised astrometry by providing high-precision measurements of the positions, motions, and parallaxes of more than a billion stars (Gaia Collaboration et al. 2018, 2021, 2023). Gaia’s precise astrometric data are particularly suited for identifying microlensing events. Belokurov & Evans (2002) estimated that approximately 25,000 stars will exhibit noticeable centroid shifts during the mission. Gaia results are published in Data Releases (DRs), which can be used to predict, find, or confirm astrometric microlensing events. Specifically, Klüter, J. et al. (2018); Klüter et al. (2022) and Su et al. (2023) used Gaia data to predict between 4,500 and 5,672 astrometric microlensing events from 2010 to 2070.
@@
Despite the high number of expected astrometric microlensing events detected by Gaia, confirmed observations remain limited to a single candidate. This candidate was identified by Jabłońska, M. et al. (2022) among the 363 photometric microlensing events in Gaia DR3 (Wyrzykowski, Ł. et al. 2023). With the upcoming Gaia Data Release 4 in 2026, which will provide the complete astrometric time-series for all stars, there is a growing requirement for software tools to filter and detect astrometric microlensing events. By not relying on photometric data, such software would allow us to detect and study a larger sample of microlensing events, thereby enhancing our understanding of lensing objects. Chen et al. (2023) developed a data analysis pipeline for detecting astrometric microlensing events caused by black holes, with an emphasis on obtaining dark matter constraints. Their approach focuses on the detection of black holes using astrometry-only data of Gaia DR4, while in our work, we generalise the approach to other types of remnants.
@@
In this paper, we present the development and application of the Gaia Astrometric Microlensing Events (GAME) Filter. This software tool is designed to identify microlensing events within Gaia’s dataset and derive the properties of the lensing objects. We tested and validated the performance of GAME Filter using mock Gaia observations across a range of magnitudes, numbers of observations, and durations of events. Additionally, we assessed the rate of false-positive detections involving binary systems misclassified as microlensing events. As it will be shown in this paper, GAME Filter successfully recovers microlensing parameters. This indicates that GAME Filter holds great potential for providing insights into microlensing events and their associated lensing objects. We note that, at present, Gaia data has been publicly released only up to DR3, which provides astrometric solutions including positions, proper motions, and parallaxes, but does not include the underlying astrometric time-series data necessary for direct microlensing detection. However, DR4, scheduled for release in 2026, will provide these detailed astrometric measurements, enabling the detection of astrometric microlensing events. Consequently, in this study, we validate GAME Filter using mock Gaia DR4 datasets.
@@
The paper is structured as follows. In Section 2, we describe the basics of astrometric microlensing, the process of generating mock Gaia observations, and the main properties of GAME filter. In Section 3, we present the results, focusing on the performance of GAME Filter under various conditions. Section 4 provides a discussion, and Section 5 our main findings.
@@ 2 Methodology
In this section, we present the main astrometric microlensing relations. We describe how we generated mock Gaia observations with astromet [^1] https://github.com/zpenoyre/astromet.py and the main functionalities of GAME filter adapted for jaxtromet [^2] https://github.com/maja-jablonska/jaxtromet.py. astromet and jaxtromet are Python packages developed to simulate astrometric tracks of single sources, microlensing events, and binary systems (e.g., Penoyre et al. 2022). astromet is optimised for parallel CPU processing, while jaxtromet leverages the Python library JAX [^3] https://github.com/google/jax to perform calculations on GPUs, enhancing computational efficiency. Simulating 50,000 mock Gaia observations requires approximately 8 hours, while filtering and identifying microlensing events takes an additional 13 hours on a CPU with 24 cores.
@@ 2.1 Astrometric microlensing
We define $u_{0}$ as the angular impact parameter — the closest approach between the lens and the source, which occurs at time $t_{0}$ — divided by $\theta_{\mathrm{E}}$. The angular separation between the source and the lens, divided by $\theta_{\mathrm{E}}$, is thus given by
$u(t)=\sqrt{\left(\frac{t-t_{0}}{t_{\mathrm{E}}}+\delta\tau_{1}\right)^{2}+\left(u_{0}+\delta\tau_{2}\right)^{2}}.$ (3)
$\delta\tau_{1}$ and $\delta\tau_{2}$ are the two components of the microlensing parallax displacement vector $(\delta\tau_{1},\delta\tau_{2})$, which is perpendicular to the line of sight to the lens and source (Gould 2004). To express this vector and describe the impact of microlensing parallax, we define a right-handed coordinate system aligned with the source position, where $\boldsymbol{\hat{n}}$ and $\boldsymbol{\hat{e}}$ are the unit vectors pointing north and east, respectively. The Sun’s position projected onto this frame is given by
$(s_{n},s_{e})=(\Delta\mathbf{s}\cdot\boldsymbol{\hat{n}},\Delta\mathbf{s}\cdot\boldsymbol{\hat{e}}),$ (4)
where $\Delta\mathbf{s}$ is the Sun’s positional offset in the geocentric frame. The displacement vector due to microlensing parallax is then expressed as
$(\delta\tau_{1},\delta\tau_{2})=(\pi_{\mathrm{E},N}s_{n}+\pi_{\mathrm{E},E}s_{e},\,-\pi_{\mathrm{E},E}s_{n}+\pi_{\mathrm{E},N}s_{e}).$ (5)
$\pi_{\mathrm{E},N}$ and $\pi_{\mathrm{E},E}$ are the north and east components of the microlensing parallax vector $\boldsymbol{\pi}_{\mathrm{E}}$. $\boldsymbol{\pi}_{\mathrm{E}}$ is a two-dimensional vector parallel to the relative proper motion of the lens-source system and can be expressed as
$\pi_{\mathrm{E}}=|\boldsymbol{\pi}_{\mathrm{E}}|=\sqrt{\pi_{\mathrm{E},N}^{2}+\pi_{\mathrm{E},E}^{2}}.$ (6)
This, in turn, allows us to determine the lens parallax using
@@
where $\varpi$ is the parallax of the source.
@@
In general, the light from the source can blend with a luminous lens, affecting the centroid deviation described by Equation 9 (Hog et al. 1995; Miyamoto & Yoshii 1995; Walker 1995). However, in our study, we focused on dark lenses and ignored any additional contribution to the light observed from sources not related to the event. In this scenario, the shift of the light centroid traces out an ellipse in the sky (e.g., Dominik & Sahu 2000). The shape and orientation of this ellipse could, in principle, provide information about the lensing parameters, including the mass and distance of the lens. By fitting the theoretical model to the observed astrometric shifts, it is possible to derive the properties of the lensing object (Belokurov & Evans 2002).
@@ 2.2 Generating mock Gaia observations
We generated Gaia mock datasets corresponding to Gaia DR4 run from 2014.5 to 2020 with astromet. To obtain the Gaia scanning laws, the scanning angles $\varphi_{\mathrm{obs}}$ as a function of the observation time $t_{\mathrm{obs}}$, we used the Gaia Observing Schedule Tool (GOST) [^4] https://gaia.esac.esa.int/gost/. We set the reference epoch at $t_{\mathrm{ref}}=2017.5$ and assumed that there is no blending in the observations (neither from the lens nor from unrelated sources). The position and motion of a source on the celestial sphere at $t_{\mathrm{ref}}$ is defined by its Right Ascension $\alpha_{0}$, Declination $\delta_{0}$, proper motion in Right Ascension $\mu_{\alpha^{*}}$, [^5] We used a “*” symbol for motions corrected for $\cos\delta$; $\mu_{\alpha^{*}}=\mu_{\alpha}\cos\delta$. proper motion in Declination $\mu_{\delta}$, and parallax $\varpi$. We generated six microlensing mock datasets. Three data sets correspond to three different baseline G-band magnitudes $G_{0}=14$, 16.5, and 19 mag, assuming the source position $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ})$. Three data sets correspond to three different source positions $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ}),\,(277.8^{\circ},-68.3^{\circ}),\,(287.9^{\circ},45.9^{\circ})$ as shown in Table 1, assuming $G_{0}=14$, which are characterized by the number of Gaia visits $N_{\mathrm{v}}=281$, 209, and 91, respectively. [^6] These $N_{\mathrm{v}}$ values correspond to regions most frequently visited by Gaia (281 visits), regions with an approximately average visit number (91 visits), and intermediate regions (209 visits). We determine $N_{\mathrm{v}}$ using the scanninglaw Python package (https://github.com/gaiaverse/scanninglaw), as this information is not available in GOST. However, since this package does not provide data for Gaia DR4, we used $N_{\mathrm{v}}$ for Gaia DR3.. We also generate one additional data set with $t_{0}\in[2010,2014]\,$yr, $G_{0}=14$, $(\alpha_{0},\delta_{0})=(6.5^{\circ},-47.3^{\circ})$, meaning that the peak of the microlensing signal was located outside of the DR4 time span. To generate mock Gaia observations of microlensing events, we used five single-source parameters and six microlensing parameters. The single source parameters are: $\alpha_{0}$, $\delta_{0}$, $\mu_{\alpha^{*}}$, $\mu_{\delta}$, and $\varpi$. The microlensing parameters are $u_{0}$, $\theta_{\mathrm{E}}$, $t_{0}$, $t_{\mathrm{E}}$, $\pi_{\mathrm{EE}}$, and $\pi_{\mathrm{EN}}$. The specific ranges of microlensing parameters that we used to generate mock data are shown in Table 1. For each mock dataset, we simulate 50,000 microlensing events.
Table 1: Microlensing parameter ranges.
@@
 [^7] Microlensing parameter ranges used to generate mock Gaia astrometric observations. We consider three different sky positions observed with Gaia, while the other parameters are generated from random uniform distributions within the specified ranges.
@@
Other types of astrometric events can produce light centroid tracks similar to those of microlensing events, which can lead to event misclassification. To assess the rate of such false-positive microlensing detections, we generated a mock Gaia dataset of 50,000 binary system events using a procedure similar to that used for microlensing events. The required parameters are five single-source parameters and nine binary system parameters. The parameters of binary systems are the orbital period $P$, the orbital semimajor axis $a$, the orbital eccentricity $e$, the ratio between masses of both stars $q$, the ratio between luminosities of both stars $l$, the polar viewing angle $\theta_{\mathrm{v}}$, the azimuthal viewing angle $\phi_{\mathrm{v}}$, the planar projection angle $\omega_{\mathrm{v}}$, and time of periapse passage $t_{\text{peri}}$ (Penoyre et al. 2022). The specific ranges of the binary system parameters used to generate mock data are gathered in Table 2.
Table 2: Binary system parameter ranges.
@@
 [^8] Binary system parameter ranges used to generate mock Gaia astrometric observations. We consider a specific $t_{\mathrm{peri}}$, while the other parameters are generated from random uniform distributions limited by the specific parameter range. The single source parameter ranges and distributions are the same as for microlensing events (see Table 1).
Table 3: Mock datasets.
@@
 [^9] Mock datasets with their respective parameters and accuracy of parameter estimation. In the first column, we list the names of all mock datasets, where we adopt the following naming convention. The prefixes “lens” and “binary” indicate a microlensing or binary mock dataset, respectively. The number following “G” refers to the baseline G-band magnitude of the source $G_{0}$. The number following “N” represents the number of Gaia visits $N_{\mathrm{v}}$ for that particular sky position. The suffixes “DR4” and “extended” denote whether the mock dataset assumes that the time of closest approach $t_{0}$ falls within the Gaia DR4 time span or extends beyond it, respectively.
In Table 3, we list all the mock datasets. We assumed a uniform distribution of single-source, microlensing, and binary parameters listed in Tables 1 and 2. This allowed us to systematically evaluate GAME Filter’s performance across a wide range of potential lensing configurations without being restricted to specific astrophysical assumptions or parameter distributions. While a more physically motivated distribution of lens and source properties (e.g., derived from Galactic models; e.g. Mao & Paczyński 1996) would be necessary for a statistical study of microlensing rates, we focused on assessing the filter’s robustness and sensitivity rather than estimating event occurrence rates.
2.3 GAME Filter
GAME Filter [^10] https://github.com/tajjankovic/GAME-Filter/ is a software tool developed to identify microlensing events in the Gaia dataset and obtain the parameters of the event that can help derive parameters and properties of the lensing object. The software reads $x_{\mathrm{obs}}$, $x_{\mathrm{err}}$, $\Delta x_{\mathrm{obs}}$, $t_{\mathrm{obs}}$, and $\varphi_{\mathrm{obs}}$ from the Gaia data files. GAME Filter then calculates $x_{\mathrm{fit}}$, which is the deviation along $\varphi_{\mathrm{obs}}$ at $t_{\mathrm{obs}}$, for specific single source and microlensing parameters. After that, the software minimises the renormalized microlensing unit weight error (MUWE)
a scalar parameter which indicates the goodness of the microlensing fit, where $N$ is the number of visits for a specific event. The goal of the minimisation is to identify parameters that bring MUWE as close as possible to its theoretical minimum value of 1. The minimisation process uses the limited memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS-B) algorithm [^11] We tested several minimisation methods and found that the L-BFGS-B method produces the most accurate results. This is mainly due to its feature of constraining individual parameters within bounds to explore the parameter space and determine the optimal single source and microlensing parameters for individual events. Appendix A provides a more detailed description of the minimisation process.
Following the minimisation process, the minimiser might stop at an incorrect local minimum, failing to find the correct solution. Consequently, we established criteria to determine when an event is recovered. These criteria are based on the value of MUWE after minimisation $\mathrm{MUWE}_{\mathrm{min}}$, L2 optimality error $L_{\mathrm{opt}}$ (see Appendix A), initial guesses, and the boundaries imposed on individual parameters. We determined the critical values of $\mathrm{MUWE}_{\mathrm{min}}$ and $L_{\mathrm{opt}}$ from their respective histograms as the values where histograms show a sharp decline. We considered an event to be recovered if the following criteria were met.
  • $1.10.9<\mathrm{MUWE}_{\mathrm{min}}<$ 1.1.
  • $L_{\mathrm{opt}}<0.015$.
  • The values of all parameters are within the boundaries imposed.
3 Results
In Table 3, we list all the mock datasets with their abbreviated names and properties. Furthermore, we show the percentages of recovered events $P_{\mathrm{rec}}$ and the percentages of events with parameter values obtained after minimisation within 10% (20%) of their true values, denoted as $P_{10}$ ($P_{20}$). These metrics provide a direct measure of GAME filter’s accuracy. Detailed discussions on these results are provided in the following subsections.
The true parameter values used to generate mock observations are: the source’s Right Ascension ($\alpha_{0}$), Declination ($\delta_{0}$), proper motion in Right Ascension ($\mu_{\alpha^{*}}$), proper motion in Declination ($\mu_{\delta}$), and parallax ($\varpi$). The microlensing parameters are: the impact parameter ($u_{0}$), the Einstein radius ($\theta_{\mathrm{E}}$), the time of the closest approach between the source and the lens ($t_{0}$), the event timescale ($t_{\mathrm{E}}$), and the microlensing parallax components ($\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$). We refer to the parameter values obtained after minimisation as $\alpha_{0,\mathrm{min}}$, $\delta_{0,\mathrm{min}}$, $\mu_{\alpha^{*},\mathrm{min}}$, $\mu_{\delta,\mathrm{min}}$, $\varpi_{\mathrm{min}}$, $u_{\mathrm{0,min}}$, $\theta_{\mathrm{E,min}}$, $t_{\mathrm{0,min}}$, $t_{\mathrm{E,min}}$, $\pi_{\mathrm{EE,min}}$, and $\pi_{\mathrm{EN,min}}$. In our study, we omit showing the plots for $\alpha_{0,\mathrm{min}}$ and $\delta_{0,\mathrm{min}}$ because we impose tight boundaries on these parameters, and they are always accurately recovered. This is motivated by the fact that the position of the observed region in the sky is always accurately known.